# Infrared Small Target Detection in Satellite Videos: A New Dataset and A Novel Recurrent Feature Refinement Framework

### v2

@@ V-C2 Pyramid Deformable Alignment
Figure 13: Ablation on the spatial modulation (SM) and frequency modulation (FM) of TSFM. We show the zoom-in target regions of image, features before and after SM, FM. The zoom-in spatial attention are also presented.
@@
Table VI shows the comparative results achieved by pyramid deformable alignment and its variants. It can be observed that deformable convolution for feature alignment is superior than optical flow-based method for 5.011 $P_{d}$ gain and 34.913 $F_{a}$ drop. This is because, deformable convolution employs implicit motion compensation in an unified step, avoiding ambiguous and duplicate results caused by wrapping. In addition, attention-based methods perform the worst, and even worse than flow-based method. This is because, the complex background clutter and variable target appearance introduce much noise for spatial weighted summation operation of attention-based methods. In addition, pyramid structure can introduce 0.853 $P_{d}$ gain, which further demonstrates the effectiveness of coarse-to-fine feature fusion. In Fig. 11, we show adjacent images, corresponding features, together with aligned features, and depict the flow (derived by RAFT [86]) between reference and aligned features. Compared with the flow without feature alignment, the flow of alignment modules are much smaller and cleaner. Among them, PDA achieves the smallest and cleanest flow results. It is demonstrated that PDA module can successfully compensate motions introduced by moving satellite.
@@ V-C3 Temporal-Spatial-Frequent Modulation
Table VII shows the comparative results achieved by temporal-spatial-frequent modulation (TSFM) and its variants. It can be observed that temporal modulation introduces 2.454 $P_{d}$ gain and 2.637 $F_{a}$ drop compared with ResUNet-w/o, which demonstrates its effectiveness for feature fusion. In Fig. 12, we visualize the flow between the reference feature and aligned feature, together with the temporal attention map. It can be observed from the visualization of flow and temporal attention that neighborhood features can be accurately aligned, and most of the background clutter are suppressed while target are correctly enhanced and propagated.
Spatial modulation (SM) and frequency modulation (FM) are used for feature enhancement. Ablation results in Table VII show substantial performance improvements introduced by SM and FM (1.706 and 1.706 gain for $P_{d}$). In addition, we visualize the spatial attention map, together with the features and corresponding zoom-in target region before and after SM and FM in Fig. 13. It can be observed that spatial attention and frequency attention effectively enhance the response of high-intensity and high-frequency regions, especially for small objects.

### v3

@@ V-C2 Pyramid Deformable Alignment
Figure 13: Ablation on the spatial modulation (SM) and frequency modulation (FM) of TSFM. We show the zoom-in target regions of image, features before and after SM, FM. The zoom-in spatial attention is also presented.
@@
Table VI shows the comparative results achieved by pyramid deformable alignment and its variants. It can be observed that deformable convolution for feature alignment is superior than optical flow-based method for 5.011 $P_{d}$ gain and 34.913 $F_{a}$ drop. This is because, deformable convolution employs implicit motion compensation in a unified step, avoiding ambiguous and duplicate results caused by wrapping. In addition, attention-based methods perform the worst, and even worse than flow-based methods. This is because, the complex background clutter and variable target appearance introduce much noise for spatial weighted summation operation of attention-based methods. In addition, pyramid structure can introduce 0.853 $P_{d}$ gain, which further demonstrates the effectiveness of coarse-to-fine feature fusion. In Fig. 11, we show adjacent images, corresponding features, together with aligned features, and depict the flow (derived by RAFT [94]) between reference and aligned features. Compared with the flow without feature alignment, the flows of alignment modules are much smaller and cleaner. Among them, PDA achieves the smallest and cleanest flow results. It is demonstrated that PDA module can successfully compensate motions introduced by moving satellite.
@@ V-C3 Temporal-Spatial-Frequent Modulation
TABLE VIII: $P_{d}$(×$\times 10^{2}$) and $F_{a}$(×$\times 10^{6}$) values achieved by ResUNet_RFR on 5 image sequences with one-pixel targets and 5 image sequences with sub-pixel targets.
One-pixel targets Sub-pixel targets Total
$P_{d}$$\uparrow$ $F_{a}$$\downarrow$ $P_{d}$$\uparrow$ $F_{a}$$\downarrow$ $P_{d}$$\uparrow$ $F_{a}$$\downarrow$
86.03 30.80 82.52 18.77 84.37 24.95
Figure 14: Qualitative results of ResUNet_RFR method on one-pixel targets (row 1) and sub-pixel targets (row 2).
Table VII shows the comparative results achieved by temporal-spatial-frequent modulation (TSFM) and its variants. It can be observed that temporal modulation introduces 2.454 $P_{d}$ gain and 2.637 $F_{a}$ drop compared with ResUNet-w/o, which demonstrates its effectiveness for feature fusion. In Fig. 12, we visualize the flow between the reference feature and aligned feature, together with the temporal attention map. It can be observed from the visualization of flow and temporal attention that neighborhood features can be accurately aligned, and most of the background clutters are suppressed while targets are correctly enhanced and propagated.
Spatial modulation (SM) and frequency modulation (FM) are used for feature enhancement. Ablation results in Table VII show substantial performance improvements introduced by SM and FM (2.176 and 1.706 gain for $P_{d}$). In addition, we visualize the spatial attention map, together with the features and corresponding zoom-in target region before and after SM and FM in Fig. 13. It can be observed that spatial attention and frequency attention effectively enhance the response of high-intensity and high-frequency regions, especially for small objects.
V-D Discussion
In this subsection, we investigate the generalization of RFR on one-pixel targets and sub-pixel targets. Specifically, we synthesized 5 image sequences with one-pixel targets and 5 image sequences with sub-pixel targets. Based on the synthetic data, we employ ResUNet_RFR as the baseline algorithm to conduct evaluation experiments. Quantitative and qualitative results in Table VIII and Fig 14 show that ResUNet_RFR can perform robust detection on one-pixel and sub-pixel targets.